Determaine the proability that next years production will exceed the production level 107.6. Mean = 212.1 and STDEV = 123.2997857.
Answer by stanbon(75887)

Determaine the proability that next years production will exceed the production level 107.6. Mean = 212.1 and STDEV = 123.2997857.
-----------------------
Find the z-score of 107.6
---
z(107.6) = (107.6-212.1)/123.3 = -0.8455
---------
P(x>107.6) = P(z>-0.8455)
Using your z-chart or your calculator to find the following:
P(z > -0.8455) = 0.8011
---------------------------------------
Cheers,
Stan H.